# Separating subsets from their images

### v1

@@ 7. $(vii)$
    $C_{2}\mathop{\mathrm{wr}}\mathrm{Alt}(n)$ endowed with the imprimitive action;
@@ 8. $(viii)$
    $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(n)$ endowed with the imprimitive action;
@@
We start with the only subgroups of $C_{2}\mathop{\mathrm{wr}}H$ (where $H$ is $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$) that induce $H$ on the blocks and for which we show that they do not meet the upper bound.
We analyze in turn the possible groups we have described in the discussion preceding this lemma. In what follows, $X$ denotes a $k$-subset of the domain, with $k\leq n$, and $\Sigma_{0},\Sigma_{1}$ and $\Sigma_{2}$ denote the sets of blocks that contain 0, 11 and $2$ points from $X$, respectively.
Let us start with those cases where $K$, the kernel of the action on the block system $\Sigma$, is trivial. The only possibility is that $G=\mathrm{Sym}(n)$, and its stabiliser is $\mathrm{Alt}(n-1)$. In particular, $\mathrm{Alt}(n)\leq G$ acts intransitively defining two orbits of size $n$, while every odd permutation, along with its action on the $n$ blocks, swaps the two elements within each block. By $n$-transitivity of the action of $G$ on the blocks, there exists an odd permutation $g$ that exchanges $\Sigma_{0}$ and $\Sigma_{2}$. (Since $n\geq 4$, if $g^{\prime}$ is an even permutation with the property, then $g$ is an odd permutation obtained by multiplying $g^{\prime}$ by a transposition that exchanges two blocks in a $\Sigma_{i}$, for some $i\in\{0,1,2\}$). By our description of $G$, it is immediate to verify that $X\cap X^{g}$ is empty. This completes the proof of $(i)$. Note that $(vi)$ contains a subgroup permutationally isomorphic to $G$, and hence it enjoys the required property.
Suppose that $K=D$ is the diagonal subspace. If $H$, the group that $G$ induces on $\Sigma$, is alternating, we can determine $g$ such that $X\cap X^{g}$ is empty using similar ideas as before. By (n2)(n-2)-transitivity of $\mathrm{Alt}(n)$, there exists a permutation $h$ that swaps the sets $\Sigma_{0}$ and $\Sigma_{2}$. If we compose $h$ with $s$, the nontrivial element of $D$ that swaps the elements within each blocks, we immediately see that $X\cap X^{hs}$ is empty. Thus, $g=hs$, and $(ii)$ is obtained. Observe that $(iii)$, $(vii)$ and $(viii)$ follows immediately as $\mathrm{Alt}(n)\times C_{2}$ is a subgroup of the permutation groups into consideration. Note that we have exhausted all the cases with $K\in\{D,\mathbb{F}_{2}\Omega\}$.
Last, we suppose that $K=A$. We need to consider two cases according to $n$ being even or odd. Observe that $C_{2}=D\leq A$ if, and only if, $n$ is even. Therefore, if $n$ is even $\mathrm{Alt}(n)\times C_{2}$ is a subgroup of the groups in question: this takes care of $(iv)$ and $(v)$ Assume that $n$ is odd and that $H$ can either be alternating or symmetric. Note that, as the extension is split, $H$ defines two orbits of size $n$. Let $X$ be one of them. Then any $g\in G$ can be uniquely written as $ha$, for some $a\in A$ and for some $h\in H$. As $X^{ha}=X^{a}$, we need to find $a$ such that $X\cap X^{a}$ is empty. But, since every $a\in A$ swaps the two elements inside an even number of blocks, this is not possible. Hence there does not exist a $g\in G$ that such that $X\cap X^{g}$ is empty. In particular, $\mathbf{m}(G)\leq n$ and $G$ does not appear in the list. This concludes the proof, as all the transitive groups of $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(n)$ that induces either $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ on $\Sigma$. ∎
Suppose that $G$ is imprimitive with two blocks of cardinality $n$. Let us denote these blocks by $B$ and $C$ and let $X,Y$ be two subsets of $B$ of equal size $k<n/2$. We will show that there exists a permutation in $G$ mapping $X$ to $Y$. Let $g$ be an arbitrary permutation that swaps the two blocks $B$ and $C$. Consider the subset
$A=X\cup\left(C\setminus Y^{g^{-1}}\right).$
Since $|A|=n<\mathbf{m}(G)$, there exists a permutation $h$ such that $A\cap A^{h}$ is empty. Observe that $X^{h}=Y^{g^{-1}}$ and hence $X^{hg}=Y$. By the arbitrariness of $X$ and $Y$, we conclude that the permutation group that $G^{B}_{B}$, induced by the action of $G_{B}$ on $B$, is $k$-homogeneous, for every $k<n/2$. Using again [13, Section 5] and [18, Theorem 9.4B], we have that the remaining possibilities for $G_{B}^{B}$ are as follows:
  1. $(e)$
    $\mathrm{Alt}(n)\leq G_{B}^{B}\leq\mathrm{Sym}(n)$ in their natural action on $n$ points;
  2. $(f)$
    $\mathrm{ASL}_{d}(q)\leq G_{B}^{B}\leq\mathrm{A}\Gamma\mathrm{L}_{d}(q)$ in their natural action on $q^{d}$ points, where
    $(d,q)\in\{(1,4),(1,5),(1,8),(2,2),(3,2)\};$
  3. $(g)$
    $\mathrm{PSL}_{d}(q)\leq H\leq\mathrm{P}\Gamma\mathrm{L}_{2}(q)$ in their natural action on $\mathrm{PG}_{1}(q)$, where $q\in\{4,5,7,8\}$;
  4. $(h)$
    $G_{B}^{B}$ is isomorphic to $M_{12}$ in its natural actions on $12$ points.
Once again, we need to consider these cases in turn. We postpone cases (f)(f)$(g)$ and  $(h)$ to the section of the proof where we collect our algorithmical considerations. In contrast, we cannot handle case $(e)$ computationally. The structure of the groups satisfying all the hypothesis we imposed can be quite wild, but can have a good understanding of some relevant subgroups for our purposes.
Assume that $n\geq 5$, and let $G$ be an imprimitive permutation group with two blocks, $B$ and $C$, of size $n$ such that $\mathrm{Alt}(n)\leq G_{B}^{B}\leq\mathrm{Sym}(n)$. We compute the socle of $G$. Preliminary, we note that $G_{(\Sigma)}=G_{B}=G_{C}$ and that $G$ is an extension of $C_{2}$ by $G_{(\Sigma)}$. Let $T$ be a minimal normal subgroup of $G_{(\Sigma)}$. Consider the projections $\pi_{B}:G_{(\Sigma)}\to G_{B}^{B}$ and $\pi_{C}:G_{(\Sigma)}\to G_{C}^{C}$. Since $T$ is minimal normal, so are $T^{\pi_{B}}$ and $T^{\pi_{C}}$. Observe that, as $G$ is faithful, the two images cannot be concurrently trivial. In particular, $\mathrm{Alt}(n)$ is a subgroup of $T$, and hence the socle of $G$ is of the form $\mathrm{Alt}(n)^{\ell}$, for some positive integer $\ell$. As $G\leq\mathrm{Sym}(n)\mathop{\mathrm{wr}}C_{2}$, we see that $\ell\leq 2$, and hence either $\ell=1$ or $\ell=2$. If $\ell=2$, it follows that
$\mathrm{Alt}(n)\mathop{\mathrm{wr}}C_{2}\leq G\leq\mathrm{Sym}(n)\mathop{\mathrm{wr}}C_{2}.$
On the other hand, suppose that $\ell=1$. Since $G$ acts transitively on $\Sigma$, there exists a permutation $s\in G$ such that $B^{s}=C$ and $s^{2}\in G_{(\Sigma)}$. In particular, depending on whether $G_{B}^{B}$ is alternating or symmetric, $G$ contains a subgroup isomorphic to either $\mathrm{Alt}(n)\langle s\rangle$ or $\mathrm{Sym}(n)\langle s\rangle$. Observe that $G_{B}^{B}$ and $G_{C}^{C}$ are permutationally isomorphic: the requested bijections are $s$ for the domains and the conjugation by $s$ for the groups. By [18, Exercise 1.6.1], there exists $f\in\mathrm{Sym}(n)$ such that every element of $G$ can be written as $(x,x^{f})$, where $x$ is a permutation of $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ in natural action. The possibilities for $G$ depend on whether the action induced on the two blocks are equivalent or not. By assuming that $n\geq 7$, and by applying, for instance, [18, Lemma 1.6B], we find that the alternating group has two (permutationally isomorphic but) not equivalent actions on $n$ points (and one is mapped to the other by applying an outer automorphism of $\mathrm{Alt}(n)$), while the symmetric group has only one. Therefore, there are three representatives for the permutational isomorphism classes for $G$.
  • If $G_{B}^{B}=\mathrm{Alt}(n)$ and $f$ is even, then $G$ is permutationally isomorphic to $\mathrm{Alt}(n)\times C_{2}$;
  • If $G_{B}^{B}=\mathrm{Alt}(n)$ and $f$ is odd, then $G$ is permutationally isomorphic to $\mathrm{Alt}(n)\rtimes\langle s\rangle$, where $s^{2}=1$ and $s$ applies a transposition while swapping the blocks;
  • If $G_{B}^{B}=\mathrm{Sym}(n)$, then $G$ is permutationally isomorphic to $\mathrm{Sym}(n)\times C_{2}$.
(We point out that, by repeating the same reasoning, we can describe the transitive permutation groups arising in case $(h)$. Indeed, as $N_{\mathrm{Sym}(12)}(M_{12})/M_{12}^{2}\cong\mathrm{Out}(M_{12})$ has order $2$, we find that $G$ is isomorphic to either $M_{12}\mathop{\mathrm{wr}}C_{2}$, to $M_{12}\times C_{2}$, or to $M_{12}\langle s\rangle$, where the $s$ swaps the blocks and the action that the subgroup $M_{12}$ induces on the two blocks are not equivalent. In our algorithm, we actually built these groups as the only three transitive subgroups of $M_{12}\mathop{\mathrm{wr}}C_{2}$, rather than using this theoretical approach.)
We can now establish which groups from $(e)$ meet the upper bound from D.
Lemma 28.
Let $n\geq 8$ be a positive integer, and let $G$ be an imprimitive group with block system $\Sigma$ that consists of two blocks of size $n$. Suppose that the action of $G$ on a block is permutationally isomorphic to $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$. Then $\mathbf{m}(G)=n+1$.
Proof.
First, we have verified that the statement is true for $n=8$ with a computer-aided calculation. Hence, during the proof, we assume that $n\geq 9$. We split the discussion in two cases, depending on the socle of $G$ being isomorphic to $\mathrm{Alt}(n)^{2}$ or to $\mathrm{Alt}(n)$.
Suppose that $\mathrm{Alt}(n)\mathop{\mathrm{wr}}C_{2}\leq G$, and let $A$ be a $k$-subset of points, for some $k\leq n$. We write $k_{B}=|A\cap B|$ and $k_{C}=|A\cap C|$. If $\max\{k_{B},k_{C}\}\leq\mathbf{m}(\mathrm{Alt}(n))$, then there exists a permutation $g\in\mathrm{Alt}(n)^{2}$ such that $A\cap A^{g}$ is empty. Otherwise, as $\mathbf{m}(\mathrm{Alt}(n))>n/2$, we can assume (by eventually renaming the blocks) that
$k_{C}=k-k_{B}<n/2.$
Let $s\in C_{2}$ be the element of $G$ swapping the two blocks. Note that
$|(A\cap B)^{s}\cap(A\cap C)|\leq k_{C}<n/2\quad\hbox{and}\quad|(A\cap C)^{s}\cap(A\cap B)|\leq k_{C}<n/2.$
Thus, by $(n/2)$-transitivity of $\mathrm{Alt}(n)$, we can choose $(h_{B},h_{C})\in\mathrm{Alt}(n)^{2}$ such that
$(A\cap B)^{sh_{C}}\cap(A\cap C)=\emptyset\quad\hbox{and}\quad(A\cap C)^{sh_{B}}\cap(A\cap B)=\emptyset.$
Hence, $g=s(h_{B},h_{C})\in G$ has the property that $A$ and $A^{g}$ are disjoint. This exhausts the first case.
@@ 4. $(h)$
Suppose that $\mathrm{Alt}(n)\langle s\rangle\leq G$ but $\mathrm{Alt}(n)\mathop{\mathrm{wr}}C_{2}$ is not a subgroup of $G$. If $s$ is central in $\mathrm{Alt}(n)\langle s\rangle$, $G$ has blocks of size two, and hence the conclusion follows from Lemma 27. Assume now that $s$ is not central in $\mathrm{Alt}(n)\langle s\rangle$. Let $A$ be a $k$-subset of points, for some $4\leq k\leq n$. We can suppose that $A$ and $A^{s}$ are not disjoint. We label the points of $B$ by $\{1_{B},2_{B},\dots,n_{B}\}$, and those of $C$ by $\{1_{C},2_{C},\dots,n_{C}\}$. By the discussion preceding this lemma, we can also assume that $1_{B}^{s}=2_{C}$, $2_{B}^{s}=1_{C}$, $1_{C}^{s}=2_{B}$, $2_{C}^{s}=1_{B}$, while, for every integer $3\leq x\leq n$, $x_{B}^{s}=x_{C}$ and $x_{C}^{s}=x_{B}$. We define the integers
@@
Finally, the pointwise stabilizer of the set $\Omega-\left\{1_{B},2_{B},1_{C},2_{C},a_{B},b_{B},a_{C},b_{C}\right\}$ contains $\mathrm{Alt}(n-4)$. The choice of a suitable $h$ so that $A$ and $A^{sth}$ are disjoint is the same as in the corresponding case in Lemma 27. (Note that our proof of the equality $\mathbf{m}(\mathrm{Alt}(n)\times C_{2})=\mathbf{m}(\mathrm{Sym}(n)\times C_{2})=n+1$ does not require $n\geq 8$, but rather the weaker $n\geq 5$: this is what forces us to take $n\geq 9$ in this proof.)
Since no more transitive groups with the required properties exist, the proof is complete. ∎
We now focus on establishing which of the listed groups meet the putative upper bound. To understand which possibility in cases (b)(b)(c)(c)(f)(f)$(g)$ and $(h)$ achieve the upper bound, we run a computer-aided calculation, whose code is available at [5]. Initially, we filtered the imprimitive permutation groups of degree up to $18$ with the property that they define a block system $\Sigma$ with $G_{B}^{B}$ and $G^{\Sigma}$ as prescribed. Then, we applied a greedy approach to establish if they achieve or not the upper bound by checking whether there exists a $n$-subset of the domain that is not self-separable. (We point out that, for this algorithm, we have not used Lemma 23 – we are unaware if filtering these groups using this criteria would speed up the computation.)
This line of attack seems to be too time consuming for case $(d)$. Running the code on the single group $C_{2}\mathop{\mathrm{wr}}M_{11}$ shows that $\mathbf{m}(C_{2}\mathop{\mathrm{wr}}M_{11})\leq 11$, and hence the putative upper bound of $12$ is not reached. In particular, by Lemma 12, this excludes that any subgroup of $C_{2}\mathop{\mathrm{wr}}M_{11}$ meets the upper bound. On the other hand, $\mathbf{m}(C_{2}\mathop{\mathrm{wr}}M_{12})=13$. We should not establish which subgroup of the wreath product is transitive on $24$ points: this task is computationally too demeaning, and this is the reason why earlier we described the transitive subgroups of $C_{2}\mathop{\mathrm{wr}}M_{12}$ that induce $M_{12}$ as their permutation group on the block system. In particular, we feed these transitive groups to the algorithm that checks whether subsets of $12$-points are self-separable. We obtain that all of them are for all the groups under consideration, with the exception of the perfect central extension of $C_{2}$ by $M_{12}$. We have thus considered all the group arising in $(d)$.
Proof of Lemma 22.
The statement for permutation groups whose degree is at most $24$ follows from the preceding discussion and the computation of [5]. The remaining possibilities has been addressed in Lemma 27 and Lemma 28. No other case need to be considered, and thus our proof is complete. ∎
@@ 8.1. Almost simple type
The proof of J is divided into several lemmas depending on the socle of the group considered.
We start by considering the action of the alternating group $\mathrm{Alt}(m)$ or the symmetric group $\mathrm{Sym}(m)$ on the set $m]\choose k$} of $k$-subsets of the set $m]=\{1,2,\ldots,m\}$. Observe that the degree of this permutation representation is, $m\choose k$}. (Lemma 29 takes care of the case $(a)$ in Table 1.)
@@
This intersection is never empty, and, if $g$ does not stabilise $\{1,2,\ldots,r\}$, $A^{g}$ and $A$ are distinct. Hence, $A$ is not self-separable for $\mathrm{Sym}(m)$. This considerations prove the upper bound of $\mathbf{m}(\mathrm{Sym}(m))$.
@@ 3. $(iii)$
Let now $G$ be a classical group whose socle is not isomorphic to $\mathrm{PSL}_{d}(q)$. We start by considering the action of $G$ on the totally isotropic or totally singular $k$-subspaces of its natural module. From here on, with the symbol $V_{\bullet}^{(k)}$ we refer to the set of $k$-dimensional subspaces of the formed space $(V,\mathbf{b})$ where $\bullet$ signals a given property, which is clear from context. For instance, in the next proof $V_{\bullet}^{(k)}$ denotes in a unified way the totally isotropic and totally singular $k$-subspaces. (Lemma 31 takes care of the cases $(e)$, $(g)$, $(i)$, $(j)$ and $(k)$ in Table 1. Moreover, not to overcomplicate our formulae, we adopt the notation that if, in a product $\prod_{i=a}^{b}x_{i}$, $a$ is greater than $b$, then the product is equal to $1$.)
@@
Let $G$ be a classical group whose socle is not linear, and consider its subspace action on totally isotropic or totally singular $k$-subspaces of its natural module $V$. Then there exists $A$ which is a non self-separable subset of the domain whose exact cardinality, as well as the cardinality of the domain, are collected in Table 3.
@@
Moreover, the expression of $|\Omega|$ and $|A|$ appearing in Table 3 appear to be to complicated to obtain an explicit asymptotic behaviour. We shall be satisfied by observing that the upper bound that we have computed can be (trivially) formulated as
$|\Omega|^{\log_{|\Omega|}|A|},$
and that many experimentation suggests that, for a fixed $k$, the exponent should be arbitrarily close to $1/2$ if we allow $d$ to grow to infinity.
@@
Let $G$ be a classical group whose socle is not linear, and consider its subspace action on nondegenerate $k$-subspaces of its natural module $V$. Then there exists $A$ which is a non self-separable subset of the domain whose exact cardinality and asymptotic size in relation with the degree are collected in Table 4.
In Table 4, the letters $B$ and $C$ replace the functions
@@
As in the proof of Lemma 31, we start by fixing a subspace $U$. Unlike before, we need to make from the start different choices of $U$ depending on the socle of $G$ and, if the bilinear form is symmetric, depending if the subspaces we are acting on are hyperbolic, parabolic or elliptic. These choices are as follows.

### v2

@@ 7. $(vii)$
    one of transitive groups of degree $10$ that can be found in the Library of Transitive Groups in [9] under the names TransitiveGroup(10,i)(10,i) for $i\in\{17,19,20,27,41,$ 42}42\}.
@@ 8. $(viii)$
    one of transitive groups of degree $12$ that can be found in the Library of Transitive Groups in [9] under the names TransitiveGroup(12,i)(12,i) for $i\in\{181,182,229,278,$ 279,297,298}279,297,298\}.
  9. $(ix)$
    one of transitive groups of degree $14$ that can be found in the Library of Transitive Groups in [9] under the names TransitiveGroup(14,i)(14,i) for $i\in\{59,60\}$.
  10. $(x)$
    one of transitive groups of degree $16$ that can be found in the Library of Transitive Groups in [9] under the names TransitiveGroup(16,i)(16,i) for $i\in\{1078,1502,1798,$ 1799,1801,1802,1882,1883,1950,1951}1799,1801,1802,1882,1883,1950,1951\}.
  11. $(xi)$
    one of transitive groups of degree $18$ that can be found in the Library of Transitive Groups in [9] under the names TransitiveGroup(18,i)(18,i) for $i\in\{937,938,979,$ 980}980\}.
Remark 38.
Observe that $\mathbf{q}(\mathrm{Alt}(m),(1,2))\cong\mathop{\mathrm{mix}}(\mathrm{Sym}(m),\mathrm{Alt}(m),{\bf 1})$ and that $\mathbf{q}(H,1)\cong H\times C_{2}$ for all $m\geq 3$. Besides these groups, the following groups appear both in Lemma 31 and Lemma 37.
  • For degree $4$, $C_{2}\times C_{2}$, $C_{4}$ or $\rm Dih(4)$, all acting transitively on $4$ points;
  • For degree $8$,
    • the groups TransitiveGroup(8,i)(8,i) for $i\in\{21,22,26,28,29,30,31\}$, which appear in Lemma 37 and are identified in Lemma 31,
    • the groups TransitiveGroup(8,17)(8,17) and TransitiveGroup(8,18)(8,18), which appear in Lemma 31 and are identified in Lemma 37
    • and the group TransitiveGroup(8,19)(8,19), which is not indentified in either Lemma 31 or Lemma 37.
@@ 11. $(xi)$
As for Lemma 31, the statement has been verified computationally for $m\leq 9$ (see [5]). Therefore, we can assume that $m\geq 10$.
Let us show that all the permutation groups $G$ of degree $2m\geq 20$ listed in Lemma 37 satisfy $\mathbf{m}(G)=m+1$. As every other group appearing is an overgroup of both $\mathrm{Alt}(m)\times C_{2}$ and $\mathbf{q}(\mathrm{Alt}(m),(1,2))$, we just need to consider these two cases. Observe that $\mathrm{Alt}(n)\times C_{2}$ has blocks of size two, and hence the conclusion follows from Lemma 31.
@@
Hence, we need to consider $\mathbf{q}(\mathrm{Alt}(m),(1,2))$. Let $A$ be a $k$-subset of points, for some $4\leq k\leq n$. We can suppose that $A$ and $A^{s}$ are not disjoint. We label the points of $B$ by $\{1_{B},2_{B},\dots,n_{B}\}$, and those of $C$ by $\{1_{C},2_{C},\dots,n_{C}\}$. We can also assume that $1_{B}^{s}=2_{C}$, $2_{B}^{s}=1_{C}$, $1_{C}^{s}=2_{B}$, $2_{C}^{s}=1_{B}$, while, for every integer $3\leq x\leq n$, $x_{B}^{s}=x_{C}$ and $x_{C}^{s}=x_{B}$. We define the integers
@@
Finally, the pointwise stabilser of the set $\Omega-\left\{1_{B},2_{B},1_{C},2_{C},a_{B},b_{B},a_{C},b_{C}\right\}$ contains $\mathrm{Alt}(n-4)$. The choice of a suitable $h$ so that $A$ and $A^{sth}$ are disjoint is the same as in the corresponding case in Lemma 31.
Let us now prove that, if $\mathbf{m}(G)=m+1$ and $2m\geq 20$, then $G$ appears in the list in Lemma 37. Suppose that $G$ admits a system of imprimitivity $\Sigma$ consisting of two blocks of size $m$. Let us denote these blocks by $B$ and $C$, and let $X,Y$ be two subsets of $B$ of equal size $k<m/2$. We will show that there exists a permutation in $G$ mapping $X$ to $Y$. Let $g$ be an arbitrary permutation that swaps the two blocks $B$ and $C$. Consider the subset
$A=X\cup\left(C\setminus Y^{g^{-1}}\right).$
Since $|A|=m<\mathbf{m}(G)$, there exists a permutation $h$ such that $A\cap A^{h}$ is empty. Observe that $X^{h}=Y^{g^{-1}}$, and hence $X^{hg}=Y$. By the arbitrariness of $X$ and $Y$, we conclude that the permutation group that $G^{B}_{B}$, induced by the action of $G_{B}$ on $B$, is $k$-homogeneous, for every $k<m/2$. Using again [13, Section 7] and [17, Theorem 9.4B] and that $m\geq 10$, we see that either $\mathrm{Alt}(n)\leq G_{B}^{B}\leq\mathrm{Sym}(n)$ or $M_{12}=G_{B}^{B}$. In the latter scenario, in view of Lemma 36, we have two candidate groups: namely, $M_{12}\mathop{\mathrm{wr}}C_{2}$ and $\mathbf{q}(M_{12},1)\cong M_{12}\times C_{2}$. (Note that 33 produces no other subgroup as ${\bf N}_{\mathrm{Sym}(12)}(M_{12})=M_{12}$.) We have verified by a computer-aided calculation that $M_{1}2\mathop{\mathrm{wr}}C_{2}$ reaches the upper bound, while $M_{12}\times C_{2}$ does not. Therefore, we are left with $\mathrm{Alt}(n)\leq G_{B}^{B}\leq\mathrm{Sym}(n)$, and the conclusion follows by Examples 35 and 36. ∎
@@ 8.1. Almost simple type
The proof of K is divided into several lemmas depending on the socle of the group considered.
We start by considering the action of the alternating group $\mathrm{Alt}(m)$ or the symmetric group $\mathrm{Sym}(m)$ on the set $m]\choose k$} of $k$-subsets of the set $m]=\{1,2,\ldots,m\}$. Observe that the degree of this permutation representation is $m\choose k$}. (Lemma 39 takes care of the case $(a)$ in Table 1.)
@@
This intersection is never empty, and, if $g$ does not stabilise $\{1,2,\ldots,r\}$, $A^{g}$ and $A$ are not equal. Hence, $A$ is not self-separable for $\mathrm{Sym}(m)$. This proves the upper bound of $\mathbf{m}(\mathrm{Sym}(m))$.
@@ 3. $(iii)$
Let now $G$ be a classical group whose socle is not isomorphic to $\mathrm{PSL}_{d}(q)$. We start by considering the action of $G$ on the totally isotropic or totally singular $k$-subspaces of its natural module. From here on, with the symbol $V_{\bullet}^{(k)}$ we refer to the set of $k$-dimensional subspaces of the formed space $(V,\mathbf{b})$ where $\bullet$ signals a given property, which is clear from context. For instance, in the next proof $V_{\bullet}^{(k)}$ denotes in a unified way the totally isotropic and totally singular $k$-subspaces. (Lemma 41 takes care of the cases $(e)$, $(g)$, $(i)$, $(j)$ and $(k)$ in Table 1. Moreover, we adopt the notation that if in a product $\prod_{i=a}^{b}x_{i}$, $a$ is greater than $b$, then the product is equal to $1$.)
@@
Let $G$ be a classical group whose socle is not linear, and consider its subspace action on totally isotropic or totally singular $k$-subspaces of its natural module $V$. Then there exists a non self-separable subset $A$ of the domain whose exact cardinality, as well as the cardinality of the domain, are collected in Table 2.
@@
Moreover, the expressions for $|\Omega|$ and $|A|$ appearing in Table 2 are too complicated to yield a clean explicit asymptotic description. Instead, we content ourselves with the observation that our upper bound can be (trivially) rewritten as
$|\Omega|^{\log_{|\Omega|}|A|}.$
By numerical experimentation, for fixed $k$, the exponent $\log_{|\Omega|}|A|$ appears to approach $1/2$ as $d\to\infty$.
@@
Let $G$ be a classical group whose socle is not linear, and consider its subspace action on nondegenerate $k$-subspaces of its natural module $V$. Then there exists a non self-separable subset $A$ of the domain whose exact cardinality and asymptotic size in relation with the degree are collected in Table 3.
In Table 3, the letters $B$ and $C$ refer to the functions
@@
As in the proof of Lemma 41, we start by fixing a subspace $U$. Unlike before, we need to make different choices of $U$ depending on the socle of $G$ and, if the bilinear form is symmetric, depending whether the subspaces we are acting on are hyperbolic, parabolic or elliptic. Our choices for $U$ are as follows.